Shift Summary for Eve shift, 09.10.99

Patricia McBride (mcbride@fnpsph)
Sat, 11 Sep 1999 00:12:54 CDT

Shift Summary for Eve shift, 09.02.99

On Shift: P. McBride, Sasha Glasov

Also present: Ed, Sydney, Hogan, JasonL

Hours of beam delivered at all intensities: 6.5
Hours of beam delivered at nominal intensity: 5.0
Hours of E832 data (special runs not included): 3.5
Hours of special runs (laser scans, muon runs...): 0.0

Accounting of lost time: Beam unstable for first hour, trigger problems,
power glich
========================================================================
Runs ended this shift:

Tapes brought to FCC this shift: none

========================================================================
Accesses during shift:

Sasha and Sydney reset ET power supply after a temperature trip caused by hall
AC problems/repair

========================================================================
Overview of shift:

AC problems resolved. Hall temperature drops to 55 degrees
Trigger problems are reolved by swapping a banana module and reseating a
blossom board.
Beam lost for almost 2 hours due to a power glitch

========================================================================
Description of Shift:

run spill time comment
---------------------------------------------------------------------------
14497 shift begins with unstable low intensity beam
and beam stop closed
people are working on the AC for the hall

Etot votage tripped; paged Sydney
Sydney arrives and resets crates (He cannot reset all
the crates from upstairs, so he and Sasha make an
access.)

124 16:58 Beam Stop out; 4e12 data -- BUT trigger problems
ytf plots look wrong

132 17:08 end run to reinitialize the trigger

14498 17:17 Run starts with 4 kfrends

5 Trigger looks better, but Jason suspects something

11-18 cycle power on downstairs repeater, kfrend
reload maps for HCC
see problems in banana crates
Ed and JasonL agree to remove xview bananas
from the readout

34 problems persist; end run;
Jason swaps module; reseats blossom board

14499 18:39 Start new run; trigger looks okay now

4 DYC full [laser crate]; kfrend

~22 19:05 Small power glitch at MI, TeV; no estimate

100 20:50 Jamie from MCR calls to say that beam is on its way

101 20:52 The beam is back -- running smoothly
Found a small mystery in sciamon plots. Paged
Elizabeth.
It turned out to be a problem with the pedestals for
SA4.

243 00:00 waiting for S.T. to arrive


========================================================================
Follow up Notes on Previous Shift

========================================================================
Items to Follow up on Next Shift

update pedestals after next run